# MaskGaussian: Adaptive 3D Gaussian Representation from Probabilistic Masks

### v1

@@ 4.1 Experimental Settings
Table 1: Quantitative evaluation of our method compared to previous work, computed over three datasets. Results marked with dagger $\dagger$ have been directly adopted from the original paper, all others were obtained in our own experiments. Our method uses $\lambda_{m}$=0.1 from 19,000 to 20,000 iterations. # GS represents the number of Gaussians in millions. The best, and second best results are highlighted.
@@
MaskGaussian 27.43 0.811 0.227 1.205 384.7 23.72 0.847 0.181 0.590 558.3 29.69 0.907 0.244 0.694 637.1
Table 2: Ablation study comparing our approach with Compact3DGS, by applying masks for the full training duration using the same mask learning rate and hyperparameter $\lambda_{m}=0.0005$. The larger $\lambda_{m}$=0.001. # GS represents the number of Gaussians in millions.
@@
Compact3DGS 27.32 0.805 0.233 1.533 23.61 0.846 0.180 0.960 29.58 0.903 0.248 1.310
Ours ($+$full mask) 27.44 0.811 0.226 1.520 23.66 0.846 0.180 0.740 29.76 0.907 0.244 0.913
Ours ($+$full mask + larger $\lambda_{m}$) 27.42 0.809 0.228 1.171 23.59 0.845 0.183 0.549 29.74 0.907 0.247 0.570
@@
In addition to the original 3DGS [14], we compare our method with two pruning methods that use hand-crafted importance scores—RadSplat[28] and LightGaussian[5]—and one method that uses learned masks to prune Gaussians, Compact3DGS[17]. Comparison with LightGaussian is listed in the Appendix. For a fair evaluation, we focus solely on these methods’ pruning components, excluding unrelated elements such as NeRF initialization, Gaussian attribute vector quantization, and Spherical Harmonics distillation.
@@
We modify the CUDA rasterization code of 3DGS[14] to integrate our masking into the $\alpha$-blending process and conduct all experiments including performance evaluation on an NVIDIA RTX 4090. All models are trained for 30,000 iterations.
Figure 3: PSNR and number of Gaussians used when applying mask over the full training duration, with $\lambda_{m}$=0.0005, larger $\lambda_{m}$=0.001. Scene: Treehill.
@@ 4.2 Experimental Results
We summarize the performance of various methods in Table 1. Please refer to the Appendix for detailed metrics for each scene. Our MaskGaussian achieves pruning ratios of 62.4%, 67.7%, and 75.3% on Mip-NeRF360[1], Tanks & Temples[15], and Deep Blending[11], respectively, while nearly preserving image quality metrics. Notably, our method achieves even higher fidelity than the original 3DGS on Deep Blending, which we owe to the regularization effect introduced by pruning that helps prevent overfitting. While RadSplat[28] also highly maintains image quality on Mip-NeRF360, it demonstrates lower fidelity than our method and uses more Gaussians on Tanks & Temples and Deep Blending. By efficiently representing the scene through Gaussian pruning, we achieve rendering speedups of 2.05$\times$, 2.19$\times$, and 3.16$\times$ across the three datasets.
@@
Table 3: GPU Memory Requirement Comparison. Mask is applied for the full duration with $\lambda_{m}$=0.0005, and larger $\lambda_{m}$=0.001.
Scenes bicycle drjohnson garden truck
3DGS 15.05GB 11.83GB 16.09GB 7.04GB
Compact3DGS 13.91GB 10.29GB 14.42GB 6.22GB
Ours 13.41GB 10.49GB 13.78GB 6.18GB
Ours ($+$ larger $\lambda_{m}$) 12.30GB 9.32GB 11.82GB 5.26GB
@@ 4.3 Ablation Study
By using the default mask learning rate and loss hyperparameter $\lambda_{m}$ from Compact3DGS, our model demonstrates a significant improvement in rendering quality while utilizing slightly fewer Gaussians on Mip-NeRF360 (row 2 in Tab. 2 left). Moreover, we use significantly fewer Gaussians on Tanks & Temples and Deep Blending and also surpass the rendering quality of Compact3DGS by a large margin (row 2 in Tab. 2 middle and right).
To further reveal the effectiveness of our model on Mip-NeRF360, we increase the $\lambda_{m}$, placing a stronger penalty on the number of Gaussians. This adjustment makes the number of used Gaussians of our model distinguishable from Compact3DGS, and demonstrates that we exceed their performance on both rendering quality and Gaussian counts on all three datasets(row 2, 3 in Tab. 2).
Furthermore, as demonstrated in Fig. 3, our method consistently utilizes fewer Gaussians than Compact3DGS throughout the entire training phase, resulting in lower GPU memory consumption as shown in Tab. 3. This reduction is already significant during the densification stages of training, where the number of Gaussians tends to increase at a lower level than Compact3DGS, even as the model’s performance, measured by PSNR, remains competitive during this phase and superior afterward. The efficiency of our approach allows us to maintain high-quality results with fewer computational resources, making it especially well-suited for deployment in resource-constrained environments where both memory and processing power may be limited.
@@
Compact3DGS uses a Straight-Through Estimator (STE) to back-propagate gradients through binary masks and produce deterministic masks using a predefined threshold, while we use Gumbel Softmax to stochastically sample from the mask distribution, and does not depend on thresholds.
@@
Table 6: Ablation study comparing three approaches of applying $\lambda_{m}$. #GS in millions.
Dataset Mip-NeRF360 Tanks&Temples Deep Blending
[…]
$\lambda_{m}$=0.0005,      0-30000 iter 27.44 0.811 0.226 1.520 23.66 0.846 0.180 0.740 29.76 0.907 0.244 0.913
$\lambda_{m}$=0.003, 15000-30000 iter 27.38 0.809 0.229 1.178 23.63 0.845 0.182 0.488 29.77 0.906 0.250 0.410
$\lambda_{m}$=0.1,  19000-20000 iter 27.43 0.811 0.227 1.205 23.72 0.847 0.181 0.590 29.69 0.907 0.244 0.694
To evaluate the two components’ performance at a similar level of Gaussian points, we increase the mask loss hyperparameter $\lambda_{m}$ to penalize mask usage, encouraging it to use fewer Gaussians. As shown in Tab. 4 row 3, Gumbel Softmax with this increased $\lambda_{m}$ achieves a better trade-off than STE.
@@
To ablate the effectiveness of the proposed masked-rasterization, we replace it by multiplying mask with opacity and compare their difference. Table 5 demonstrates that directly multiplying the mask with opacity results in a significant reduction in the number of Gaussians, leading to a noticeable deterioration in rendering quality. We explain this as follows: in the sampling scenario, multiplying the mask with opacity prevents gradients from propagating to unsampled Gaussians, causing them to fall behind in the optimization process. As a result, even if these Gaussians are sampled again later, their contribution remains misaligned with those that have already been optimized with the scene, leading to a “death spiral” where they are eventually discarded.
We restore the number of Gaussian points in the mask-opacity multiplication approach by relaxing the penalization parameter $\lambda_{m}$, while applying a stronger penalization with a larger $\lambda_{m}$ for masked-rasterization. As shown in rows 1 and row 4 of Table 5, masked-rasterization significantly outperforms the alternative approach in terms of both PSNR and the number of points, resulting in more effective pruning, and validating the usefulness of receiving gradients for unsampled Gaussians.
Comparison with RadSplat Light.
In addition to the default model, RadSplat [28] offers a lightweight version by applying a higher pruning threshold for its importance scores, resulting in a significantly reduced final number of Gaussians, comparable to the number of initial SFM points. To align with this aggressive pruning, we apply masks after densification and increase $\lambda_{m}$ values accordingly.
@@
Fig. 4 shows that even under extreme pruning conditions with a number of Gaussians similar to the initial SFM points, our method more effectively preserves the rendering quality, making it more adaptable to resource-constrained settings.
Comparison with different $\lambda_{m}$.
The penalization hyperparameter $\lambda_{m}$ controls the degree of pruning. And we ablate three different ways of applying it:
  1. 1.
    $\lambda_{m}$=0.0005 during the entire training phase. This is the way used in Compact3DGS [17].
  2. 2.
    $\lambda_{m}$=0.003 during 15,000 to 30,000 iterations. This introduces penalization on masks only after the densification.
  3. 3.
    $\lambda_{m}$=0.1 during 19,000 to 20,000 iterations. This quickly prunes Gaussians in 1,000 iterations after the scene is well constructed.
We present the results in Tab. 6. In Deep Blending, the second method achieves the best PSNR with the lowest number of Gaussians, while on Mip-NeRF360 and Tanks&Temples the third method is overall better. We observe that performance varies across scenes, making it challenging to select a single strategy that excels in all scenarios. Consequently, we leave the selection of $\lambda_{m}$ as an avenue for future work.
@@ 6 Computing Gradients of Masks
7 Comparison with LightGaussian
LightGaussian [5] is a method developed to compress an already trained 3D Gaussian Splatting (3DGS) model through additional training steps. To ensure a fair comparison with out method under the same training schedule, we train LightGaussian from scratch and apply Gaussian pruning at the 20,000th iteration [^1] The configuration is used in LightGaussian’s official implementation. . Our model apples $\lambda_{m}$= 0.1 during 19,000 to 20,000 iterations and has $\lambda_{m}$= 0 elsewhere.

### v2

@@ 4.1 Experimental Settings
Table 1: Quantitative evaluation of our method compared to previous work, computed over three datasets. Results marked with dagger $\dagger$ have been directly adopted from the original paper, all others were obtained in our own experiments. MaskGaussian uses $\lambda_{m}$=0.1 from 19,000 to 20,000 iterations. # GS represents the number of Gaussians in millions. The best, and second best results are highlighted.
@@
MaskGaussian-$\alpha$ 27.43 0.811 0.227 1.205 384.7 23.72 0.847 0.181 0.590 558.3 29.69 0.907 0.244 0.694 637.1
Table 2: Ablation study: comparing our method with Compact3DGS. Ours-$\beta$ has the same training setting as Compact3DGS, and Ours-$\gamma$ has stronger penalty on the number of Gaussians. Unit: #GS (Million), time (mm:ss), size (MB).
@@
Compact3DGS 27.32 0.805 0.233 1.533 19:19 362.7 23.61 0.846 0.180 0.960 10:22 228.3 29.58 0.903 0.248 1.310 16:23 311.1
Ours-$\beta$ 27.44 0.811 0.226 1.520 19:44 362.5 23.66 0.846 0.180 0.740 09:52 175.4 29.76 0.907 0.244 0.913 15:47 215.2
Ours-$\gamma$ 27.42 0.809 0.228 1.171 17:52 274.2 23.59 0.845 0.183 0.549 08:54 129.8 29.74 0.907 0.247 0.570 14:08 134.9
@@
In addition to the original 3DGS [14], we compare our method with two pruning methods that use hand-crafted importance scores—RadSplat [28] and LightGaussian [5]—and one method that uses learned masks to prune Gaussians, Compact3DGS [17]. For a fair evaluation, we focus solely on these methods’ pruning components, excluding unrelated elements such as NeRF initialization, Gaussian attribute vector quantization, and Spherical Harmonics distillation.
@@
We modify the CUDA rasterization code of 3DGS [14] to integrate our masking into the $\alpha$-blending process and conduct all experiments including performance evaluation on an NVIDIA RTX 4090. All models are trained for 30,000 iterations, unless stated otherwise. We use 3 settings for hyperparameter $\lambda_{m}$, Ours-$\alpha$: $\lambda_{m}$=0.1 from 19,000 to 20,000 iterations; Ours-$\beta$: $\lambda_{m}$=0.0005 from 0 to 30,000 iterations; Ours-$\gamma$: $\lambda_{m}$=0.001 from 0 to 30,000 iterations.
Figure 3: Test PSNR and number of Gaussians used when applying mask over the entire training duration. Ours-$\beta$ has the same $\lambda_{m}$ as Compact3DGS. Scene: Treehill.
@@ 4.2 Experimental Results
We summarize the performance of various methods in Tab. 1. Please refer to the Appendix for detailed metrics for each scene. Our MaskGaussian achieves pruning ratios of 62.4%, 67.7%, and 75.3% on Mip-NeRF360 [1], Tanks & Temples [15], and Deep Blending [11], respectively, while nearly preserving image quality metrics. Notably, our method achieves even higher fidelity than the original 3DGS on Deep Blending, which we attribute to the regularization effect of pruning, helping to mitigate overfitting. By efficiently representing the scene through Gaussian pruning, we achieve rendering speedups of 2.05$\times$, 2.19$\times$, and 3.16$\times$ across the three datasets.
@@
Post-Training Results.
Our method is designed for training from scratch but can also be applied to an existing 3DGS for fine-tuning and Gaussian pruning. We compare our model with LightGaussian [5], finetuning both for 5,000 iterations on a 3DGS trained for 30,000 iterations. As shown in Tab. 5, our model achieves superior performance in the post-training setting, particularly in the LPIPS metric.
Integration with Taming-3DGS.
We demonstrate that our method can be seamlessly integrated with state-of-the-art 3DGS optimization frameworks, such as Taming-3DGS [22]. As shown in Tab. 4, using masks to prune Gaussians is a more effective approach to improving efficiency compared to directly controlling the target number of Gaussians in Taming-3DGS.
Table 3: Peak GPU Memory Requirement. Our method can prune more Gaussians and preserve equal or better quality, thus saving the most GPU memory.
Scenes counter drjohnson room truck flowers
3DGS 5.41GB 9.03GB 5.55GB 5.70GB 7.79GB
RadSplat 4.77GB 8.98GB 4.76GB 5.18GB 7.38GB
Compact3DGS 4.99GB 7.40GB 5.16GB 4.90GB 6.30GB
Ours-$\beta$ 3.94GB 7.99GB 3.84GB 4.48GB 6.26GB
Ours-$\gamma$ 3.53GB 6.45GB 3.65GB 4.41GB 5.94GB
Table 4: Integration with Taming-3DGS. “Budget” refers to Taming-3DGS’s hyperparameter that controls the target number of primitives. Scene: Garden. The difference in training time is negligible compared to the reference 3DGS and is therefore not highlighted.
Target Budget (M) 0.7 2.1 3.5 4.9
Taming-3DGS PSNR 26.736 27.350 27.481 27.617
##\##GS (M) 0.700 2.100 3.500 4.873
training time 04:47 07:28 10:37 13:47
Taming-3DGS + Compact3DGS PSNR 26.694 27.318 27.477 27.551
##\##GS (M) 0.595 1.623 2.440 3.207
training time 04:51 07:25 10:06 12:57
Taming-3DGS + Ours PSNR 26.717 27.391 27.522 27.616
##\##GS (M) 0.583 1.581 2.401 3.179
training time 4:58 7:31 10:14 13:10
3DGS (as ref.) PSNR: 27.38 ##\##GS (M): 5.81 training time: 32:32
Table 5: Post-Training results. The baseline is a normal 3DGS model trained for 30000 iterations, and both LightGaussian and MaskGaussian finetune 5000 iterations on it. #GS is in millions.
Dataset Mip-NeRF360 Tanks&Temples Deep Blending
[…]
Baseline 3DGS 27.50 0.811 0.224 3.188 23.62 0.847 0.176 1.841 29.54 0.904 0.244 2.820
LightGaussian 27.50 0.809 0.232 1.084 23.75 0.843 0.187 0.626 29.55 0.902 0.250 0.959
MaskGaussian 27.49 0.811 0.228 1.052 23.75 0.847 0.178 0.579 29.71 0.905 0.246 0.667
@@ 4.3 Ablation Study
By using the default mask learning rate and loss hyperparameter $\lambda_{m}$ from Compact3DGS, our model-$\beta$ demonstrates a significant improvement in rendering quality while utilizing slightly fewer Gaussians on Mip-NeRF360 (row 2 in Tab. 2 left). Moreover, we use significantly fewer Gaussians on Tanks & Temples and Deep Blending and also surpass the rendering quality of Compact3DGS by a large margin (row 2 in Tab. 2 middle and right).
To further reveal the effectiveness of our model on Mip-NeRF360, we increase the $\lambda_{m}$ to place a stronger penalty on the number of Gaussians and derive model-$\gamma$. This adjustment makes the number of used Gaussians of our model distinguishable from Compact3DGS, and demonstrates that we exceed their performance on both rendering quality and Gaussian counts on all three datasets(row 2, 3 in Tab. 2).
Furthermore, as demonstrated in Fig. 3, our method consistently utilizes fewer Gaussians than Compact3DGS throughout the entire training phase, resulting in lower GPU memory consumption as shown in Tab. 3. This reduction is already significant during the densification stages of training, where the number of Gaussians tends to increase at a lower level than Compact3DGS, even as the model’s performance, measured by PSNR, remains competitive during this phase and superior afterward.
@@
Compact3DGS uses a Straight-Through Estimator (STE) to back-propagate gradients through binary masks and produce deterministic masks using a predefined threshold, while we use Gumbel Softmax to stochastically sample from the mask distribution.
@@
To evaluate the two components’ performance at a similar number of primitives, we increase $\lambda_{m}$ to encourage fewer Gaussians. As shown in Tab. 6 row 3, Gumbel Softmax with increased $\lambda_{m}$ achieves a better trade-off than STE.
@@
To ablate the effectiveness of the proposed masked-rasterization, we replace it by multiplying mask with opacity and compare their difference. Tab. 7 demonstrates that directly multiplying the mask with opacity results in a significant reduction in the number of Gaussians, leading to a noticeable deterioration in rendering quality. We explain this as follows: in the sampling scenario, multiplying the mask with opacity prevents gradients from propagating to unsampled Gaussians, causing them to fall behind in the optimization process. As a result, even if these Gaussians are sampled again later, their contribution remains misaligned with those that have already been optimized with the scene, leading to a “death spiral” where they are eventually discarded.
We restore the number of Gaussian points in the mask-opacity multiplication approach by relaxing the penalization parameter $\lambda_{m}$, while applying a stronger penalization with a larger $\lambda_{m}$ for masked-rasterization. As shown in rows 1 and row 4 of Tab. 7, masked-rasterization significantly outperforms the alternative approach in terms of both PSNR and the number of points, resulting in more effective pruning, and validating the usefulness of receiving gradients for unsampled Gaussians.
Comparison with RadSplat-Light.
RadSplat [28] offers a lightweight version by increasing pruning threshold. We increase mask penalty $\lambda_{m}$ to match its pruning extent. Fig. 4 shows that even under extreme pruning conditions with a number of Gaussians similar to the initial SFM points, our method more effectively preserves the rendering quality, making it more adaptable to resource-constrained settings.
@@ 6 Computing Gradients of Masks
7 Comparison with LightGaussian (training from scratch)
LightGaussian [5] is a method developed to compress an already trained 3D Gaussian Splatting (3DGS) model through additional training steps. Besides the post-training setting, we compare with LightGaussian by training from scratch and apply Gaussian pruning at the 20,000th iteration [^1] This configuration follows the setup used in LightGaussian’s official implementation. . Our model apples $\lambda_{m}$= 0.1 during 19,000 to 20,000 iterations.